OpenStudy (anonymous):

Which set of side lengths defines a right triangle? A. 4, 8, 9 B. 5, 9, 10 C. 6, 10, 13 D. 7, 24, 25

OpenStudy (harley1312):

a right angle is 90 degrees so I'm guessing which ever one adds up to 90 is the correct answer.

OpenStudy (anonymous):

C^2=a^2+b^2

OpenStudy (harley1312):

sorry if this is incorrect but I think it's D.

OpenStudy (anonymous):

9^2=4^2+8^2 ? 81=16+64

OpenStudy (anonymous):

so A is not

OpenStudy (anonymous):

It's D, 49+576=625 and 25^2 is 625. Thanks for the help!

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

yes D is right